OpenStudy (elifbestegul):

Find k given that 3k + 1, k and -3 are consecutive terms of an arithmetic sequence.

Directrix (directrix):

3k +1, k, -3 Three consecutive terms of an arithmetic sequence. Find k. Did I write the problem correctly?

OpenStudy (elifbestegul):

Yes, you are right.

Directrix (directrix):

Let x be the common difference of the sequence. Then, 3k+1 + x = k and k + x = -3 ------------ 2k + x = -1 k + x = -3 ------------ Subtract to add out the x terms k = -1 + 3 = 2

Directrix (directrix):

3k +1, k, -3 when 2 is substituted for k becomes: 7, 2, -3 The common difference is 5 7-5 = 2 2 - 5 = -3 So k =2 does work. Questions: @elifbestegul

OpenStudy (elifbestegul):

Thank you. @directrix

Directrix (directrix):

You are welcome.